OpenStudy (anonymous):

If $1000 is invested at 5% compound monthly for 4 years the compound amount is given.by A=1000(1.005)^48. Given that log find log A

OpenStudy (amistre64):

log both sides ....

OpenStudy (anonymous):

How

OpenStudy (anonymous):

?

OpenStudy (amistre64):

A=1000(1.005)^48 log(A) = log (1000(1.005)^48) and work whatever log properties you can remember

OpenStudy (amistre64):

1000 = 10*10*10 log10 = 1 not too sure if the 1.005 is correct, but ill assume a certain degree of intelligence on it :)

OpenStudy (anonymous):

It is

OpenStudy (anonymous):

I still don't understand how to.get the answer

OpenStudy (amistre64):

do you know your log properties?

OpenStudy (tkhunny):

If there are 10 months to a year, then 0.005 is correct. As it is the case that there are 12 months in a year, you may wish to go with 0.05 / 12 = 0.004166667

OpenStudy (amistre64):

the basics are: log(ab) = loga + logb log(a/b) = loga - logb log(a^c) = c loga

OpenStudy (anonymous):

Not really

OpenStudy (amistre64):

and if we assume log is base 10 ... log(10) = 1

OpenStudy (anonymous):

Yea

OpenStudy (amistre64):

then its just a matter of applying those properties log(A) = log (1000(1.005)^48) = log (1000) + log(1.005)^48 = log (10^3) + log(1.005)^48 = 3 log (10) + 48 log(1.005) = 3 + 48 log(1.005)

OpenStudy (amistre64):

there is no simple way i can see to hand up that log 1.005

OpenStudy (anonymous):

Now.how.do.we getvA?

OpenStudy (amistre64):

you would under the log by base 10 ing both sides ....

OpenStudy (anonymous):

Example?

OpenStudy (amistre64):

\[log(A)=a+log(b)\] \[\Large 10^{log(A)}=10^{a+log(b)}\] \[\Large A=10^{a+log(b)}\] \[\Large A=10^{a}~10^{log(b)}\]

OpenStudy (amistre64):

i see one more move :) \[\Large A=10^{a}~10^{log(b)}\] \[\Large A=10^{a}~b\]

OpenStudy (amistre64):

of course this just gets back to: A = 1000 (1.005)^48

OpenStudy (anonymous):

So what would the answer be

OpenStudy (amistre64):

im not sure that youve made your objective clear enough to determine that .... are you simply trying to determine the value of A? if so, what was the need to go thru all that log work in the first place?